# Shahriar Noroozizadeh

### v1

@@ 2 Experiments: Implicit in-weights reasoning is learned
We define an in-weights version of the above task (Fig.˜3), where all examples are generated from a fixed graph $\mathcal{G}$ (of degree $d$ and path length $\ell$), rather than a fresh graph per example. Here, the model is made to memorize the full graph in its weights through edge-memorization examples, where the input is some node $v$, and the next-token target is an adjacent node $v^{\prime}$. Next, path-finding examples are generated from the same graph by picking as input a random leaf node ${v}_{\tt{leaf}}$ (a single token) and as target, the unique root-goal path (a sequence from ${v}_{\tt{root}}$ to ${v}_{\tt{leaf}}$). The model is trained for path-finding on a subset of such leaves from our fixed graph, and is tested on the remaining leaves. (We defer results on some harder variants of this topology to §C.2.)
@@
Figure 3: Overview of our in-weights path-star task. All examples are derived from a fixed path-star graph. Training involves two types of examples: (i) edge memorization examples (ii) path-finding examples, where the prefix is some leaf, and the target is the full path. Test examples are path examples corresponding to a held-out set of leaves.
Prior positive results of implicit in-weights reasoning are on small scales of 200 or fewer entities [76, 43] or on 2-hop tasks [158, 172]. Our first result below sharpens this finding: in a much larger-scale, much larger-hop path-star task that is adversarially constructed, the model learns implicit reasoning successfully.
@@
Setup. We use a from-scratch, decoder-only Transformer (GPT-mid) [125]; we corroborate all our findings on Mamba in §C.1. For the most stable results, we interleave the edge-memorization and path-finding examples during training as in Fig.˜3 and also use pause tokens [52]. We found it is important to provide both forward and reverse edges for edge-memorization in order to dodge the reversal curse [123, 9], but this may not be necessary for smaller graphs; see §D.1. Note that this reverse augmentation is not given for the path-finding examples. All details about the formatting and the hyperparameters are in §B, followed by additional analyses in §D.
(Success of implicit in-weights reasoning) On in-weights path-star graphs of as many as $5\times 10^{4}$ nodes, trained on $75\%$ of the total $10^{4}$ paths, both the Transformer and Mamba are able to predict unseen paths when conditioned on held-out leaves with as much as 100% accuracy (see left plots of Figs.˜4 and 8). Similar positive results for some harder graph topologies are in §C.2.
Shortly, we will isolate an even stronger instance of implicit reasoning from this task for analysis. To get there, let us scrutinize where the argument of B&N’24 may go differently in the in-weights setting for the model to succeed in Observation 2.1.
2.1.1 Where does the path-star-failure argument go wrong in-weights?
(Model may experience future-token gradients) The in-weights path-star task is solved (Observation 2.1) because Clever Hans cheats are not learned quickly enough, allowing future-token gradients to persist, teaching the model to find the right-to-left solution.
Figure 4: Success of Transformer in in-weights path-star task. (left) A next-token-trained Transformer achieves perfect or highly non-trivial accuracy on large path-star graphs $\mathcal{G}_{d,\ell}$ (Section˜2.1). (middle) Learning order of tokens. The tokens of a path are not learned in the reverse order i.e., the model does not learn the right-to-left solution. Thus, gradients from the future tokens are not critical for success (Section˜2.1.1). (right) Success of hardest-token-only task. In fact, the hardest token (the first) given the leaf is learned in isolation to non-trivial accuracy (Section˜2.1.1). Success of this $\ell$-fold composition task is hard to explain within the associative memory view (§2.2). Analogous plots for Mamba are in Fig.˜8. We invite the reader to contrast these in-weights task results with the in-context task ones in Fig.˜5.
Figure 5: Failure of Transformer in in-context path-star task. We report the failure of next-token Transformers in the in-context version of the path-star task, reproducing results from B&N’24. (left) Full path accuracy remains at chance level across different small graph sizes. (Middle) Learning order of tokens with teacherless (multi-token-trained) objective shows a clear right-to-left learning cascade. (right) The hardest (first) token given the leaf fails to be learned in isolation, contrasting sharply with in-weights success shown in Fig.˜4.
(No reverse-learning cascade) The target tokens in the in-weights path-star task are learned in no particular order by the Transformer and Mamba models (see middle plots of Figs.˜4 and 8 in contrast with Fig.˜5; or see Fig.˜24 for side-by-side comparison).
The above observation weakens Hypothesis 2.1.1, encouraging us to search for another one. A second possibility is that the second stage of the failure of B&N’24 does not trouble us in the in-weights task. Recall that in this stage, the first token is to be learned in isolation without future-token gradients. But this is an $\ell$-fold composition task, that is theoretically well-understood to be computationally hard (as we elaborate later). Perhaps, that is not the case for our in-weights task:
Testing Hypothesis 2.1.1 is easy: we train the model simply on the first token loss, instead of the loss over the full path sequence. We find that this task is trivial here, affirming Hypothesis 2.1.1, and isolating a much stronger and cleaner instance of implicit in-weights reasoning, one that shortly leads us to our main insight.
@@ 2.1.1 Where does the path-star-failure argument go wrong in-weights?
(Hardest, first token is learned in isolation) In the in-weights path-star task with edge-memorization and only first-token-training examples, both the Transformer and Mamba models learn the first token in isolation—without any intermediate supervision from other nodes in the path (Figs.˜4 and 8 (right)).
@@ 2.2 The contradiction behind learning the hardest token
The fact that the hardest (first) token is learned (as in Observation 2.1.1), we argue, is difficult to square with the abstraction that parametric memory strictly stores local associations. Concretely, in this abstraction, the first node requires composing a local associative recall function $\ell$-many times: recall the predecessor of ${v}_{\tt{leaf}}$ from the weights, namely $v_{\ell-1},$ then the predecessor of that, $v_{\ell-2}$, and so on. In short, one can write $v_{1}=\texttt{Predec}^{\ell}({v}_{\tt{leaf}})$. This recall function takes the shape of a matrix operation as described below:
(Local associative parametric memory) The model has memorized the edges of $\mathcal{G}$ such that for an edge $(u,v)$, $u\in\arg\max_{w}({\bf{\Phi}}(w){\bf{W}}_{\tt assoc}{\bf{\Phi}}(v))$ where intuitively ${\bf{W}}_{\tt assoc}$ encodes the associations, whereas the embedding $\bf{\Phi}$} is arbitrary and by itself encodes no associations in graph $\mathcal{G}$. For convenience, $\bf{\Phi}$} may be abstracted as orthogonal embeddings [161, 46, 70, 183, 23] or random embeddings [111, 22, 17, 160].
@@
With this structure however, the first-token $\ell$-fold composition task should intuitively require $\Omega(\exp(\ell))$ compute to learn—just like the first token was demonstrably hard-to-learn in the in-context path-star task [B&N’24]. Specifically, learning compositions with gradient-based methods is known to be empirically hard [141, 2, 55, 49, 1, 2] and proven to be theoretically hard [97, 28, 119, 141, 164, 144] (in a certain sense [^1] Such hardness results typically only show that some worst-case function in a class of compositional functions is hard-to-learn; no proclamations are made about how hard it is to learn a fixed function we care about. Indeed, with contrived initial conditions, a singleton function class can be provably learned [4, 3, 107]. Yet, in practice, learning a fixed function is hard, proving which has been an open question—until the recent negative results of Abbe et al. [6], Shoshani and Shamir [144]. These show how the (fixed) full parity function is hard to learn with gradient descent which, through a reduction, proves that in-context composition tasks are hard to learn [65]. We leave it as an open question to link our in-weights composition task to one of these hardness results. ). One way to intuit this is to notice that the optimization task is a search for a needle in the haystack: there is an $\exp(\ell)$ space of possible discrete compositions, and all but the correct one have an equally miserable loss value; with the loss terrain rendered flat and the gradients uninformative, the learner is forced to sift through the vast hypothesis space to find a needle. As a preliminary corroboration, in §E.1, we empirically find that models fail at this task when the embeddings are frozen.
@@
This barrier could be surmounted if, rather than providing supervision from only the end output of the composition, there was supervision from each hop, which would cast a graceful loss landscape. This could mean providing a data curriculum [136, 164, 171, 5, 32] or providing chain-of-thought supervision [165, 112] or reducing compositionality through various means. Indeed, prior positive results on implicit in-weights reasoning involve one such aid or the other. First, Khona et al. [76] have paths of varying lengths (see their Fig 11), which provides an implicit curriculum; they also provide full path supervision. Furthermore, their test and train paths overlap, reducing compositionality, potentially allowing the model to stitch substrings it has seen (an ability demonstrated as possible [161]). Other results [158, 172] are on 2-hop tasks, implying a friendly exponent. Our setting, by design, offers none of these aids. Our paths have fixed lengths (hence, no implicit curriculum), are disjoint (so, no test-train overlaps), and require as much as a $10$-fold composition (so, a daunting exponent). What we isolate, therefore, is a stronger and more sterile instance of implicit in-weights reasoning, one that is cleanly inconsistent within the associative view of parametric memory.
@@ 2.3 Two competing data structures for parametric memory
The paradox begins to resolve with observations in Khona et al. [76], Ye et al. [172] that we will flesh out in our setting. Both find that the nodes of their (smaller) graphs are embedded in a way that reflects a notion of distance, aiding their respective implicit reasoning tasks. Before extending this to our much larger graphs, we point out that this observation has profound implications. First, this presents an altogether different view of parametric memory itself: a memory of atomic facts that does not take shape as an associative lookup (over arbitrary embeddings) but as a geometry of highly-organized embeddings. Next, importantly, whereas an associative memory only makes local information accessible, a geometric memory readily betrays global multi-hop relationships—even when trained only on local associations, as we establish shortly.
@@
Figure 6: Evidence of global geometry of Transformer in path-star task. (a) In the heatmap, entry $(i,j)$ is the cosine distance between the leaf embedding of path $i$ (row) and the first-hop embedding of path $j$ (col). The clear diagonal line implies that embeddings within each path are more aligned, reflecting global structure. (b) UMAP projection of token embeddings where each point is a node embedding; color indicates path identity. Different paths form separated clusters (see Fig.˜18 for clearer image). (c) Heatmap for a model trained only on edge-memorization still reveals a level of geometry, although weaker than in (a). Analogous plots for Mamba are in Fig.˜9
(Global geometric parametric memory) The model has memorized the edges of $\mathcal{G}$ with embeddings ${\bf{\Phi}}_{\tt geom}$ such that for an edge $(u,v)$, $u\in\arg\max_{w}({\bf{\Phi}}_{\tt geom}(w)\cdot{\bf{\Phi}}_{\tt geom}(v))$ but furthermore, for any non-adjacent $(u,v)$, ${\bf{\Phi}}_{\tt geom}(u)\cdot{\bf{\Phi}}_{\tt geom}(v)$ reflects some (model’s own) notion of global closeness in the graph.
@@
The resolution. One can view the two parametric memories of Sections˜2.2 and 2.3 as two competing data structures, both representable by a deep sequence model, each yielding its own learning complexity of the hardest token (much like how a heap or an array would yield different search complexities). The local associative memory is analogous to a linked list: for the first token, this would incur a (benign) $\ell$-hop lookahead step but learning this incurs an exponential cost. The geometric memory is powerful in that it reduces this learning complexity all the way to $\Theta(1)$. For instance, we may expect that the embeddings of all nodes in path $i$ are clustered tightly around a unique path vector $\mathbf{z}_{i}$. In this data structure, learning the first token given the leaf node is a one-hop task: simply find a one-hop neighbor of the central node that is best-aligned with ${\bf{\Phi}}_{\tt geom}({v}_{\tt{leaf}}^{(i)})$. Such a structure indeed materializes in our massive graphs:
@@
(Evidence of global geometry) For our large path-star graphs, the (token) embeddings of the leaf and first node of a path are clustered closer to each other than those of other paths in both the Transformer and Mamba (Fig.˜6 and Fig.˜9).
@@
We corroborate similar geometries in a variety of small graphs in Fig.˜1 and Section˜C.3 through more direct visualizations. While these geometries help make sense of why implicit reasoning succeeds, it also leaves us with foundational questions, which we lay out in the next section.
3 The emergence of global geometric memory is not easy to explain
During training, the two types of parametric memory—two equally valid ways to fit the training data—must compete with each other. Depending on the reader’s prior, the rise of the geometric memory may seem familiar especially in light of well-known geometries of high-level concepts and features [36, 117, 102]. While this impression is valid in part, this section carefully isolates aspects that are unexpected. The nuance is that, unlike known phenomena, what is observed here is in a memorization task that lacks statistical redundancies. This yields the insight that the geometry cannot be easily explained by well-understood learning pressures, be it from the architecture, the optimizer, or the supervision.
3.1 Capacity pressures do not explain geometric memory.
In theories of generalization [110, 177], representations arise from pressures to represent the data as succinctly as possible, precluding a brute-force lookup. These capacity pressures are either explicit (e.g., the architecture or the regularizers) or implicit (e.g., biases of the gradient descent optimizer). For instance, a word may co-occur with thousands of similar contexts in training. A lookup of those specific contexts is too cumbersome. But fortunately, much redundancy exists in the data. Such redundancies are compressed away by the learner. From this arises a succinct, elegant geometry of high-level embeddings—such as the ones witnessed in neural word embedding models [101]. Perhaps, a parallel logic explains our observations.
(Explicit or implicit capacity pressure) Associative memory is either explicitly impossible to represent—due to the parameter count or the design of the architecture, especially the embedding bottleneck—or is implicitly less preferred by gradient descent. This is because geometric storage is a more succinct representation of the data than associative storage.
Indeed, at first glance, associative storage seems too verbose. All vertices need to be embedded near-orthogonally—demanding a massive embedding dimension that scales with vertex count $n\approx 10^{4}$ in our path-star graphs—followed by a massive $n\times n$ associative matrix ${\bf{W}}_{\tt assoc}$ (representing the full graph adjacency matrix). In contrast, a geometric embedding seems succinct. Each of the $d$ paths can be arranged along a unique dimension (e.g., Fig.˜1 top right), requiring only $d$-many dimensions in total (where $d\ll n$), supporting the logic of Section˜3.1.
This logic can be dismantled as follows. First, we demonstrate that the learner can learn the associative model, thus questioning the role of explicit pressures. Positive theoretical results in Nichani et al. [111] (see their Theorem 2) roughly prove that, with $m$ embedding dimensions (all frozen), the $m^{2}$ parameters in ${\bf{W}}_{\tt assoc}$ can store $m^{2}$ many associations. Our models have $m\approx 400$ which then appears reasonable to represent our large graphs associatively. However, this expressivity result is not a proper refutation: the results of Nichani et al. [111] cannot be rigorously imported to us [^2] Nichani et al. [111] assume one-to-one associations, but our graphs involve many-to-many associations which are intuitively harder to represent., and besides, there may be yet other explicit pressures that discourage associative memory during optimization (e.g., normalizations, weight decay). As a definitive refutation, we make an empirical demonstration: there are optimization settings where geometric memory naturally arises, but in those very settings, an associative memory can be learned with the embeddings frozen and all else unchanged. Thus, an associative memory is artificially possible where a geometric memory naturally arose.
(Associative memory can be artificially learned with our architectures) On various tiny graphs and various sequence models (Transformers, Mamba, neural networks) (Fig.˜1 and §C.3), a geometric memory arises naturally even though the same setup learns to represent the data with the embeddings frozen. This representation is purely associative since only one trainable layer exists (see §B.2.2).
Thus explicit capacity pressures do not force a geometric memory in place. But perhaps, a geometry still arises due to implicit pressures from gradient descent to represent succinctly. To refute this, we refute the underlying premise itself: contrary to intuition, a geometric storage is not necessarily more succinct than an associative storage. To see why, we first clarify that unlike a typical natural language task, our setting must be recognized as a memorization task (in the sense of Feldman [38]): the existence of an edge cannot be statistically surmised from the rest of the training set. In other words, while typical word embedding tasks contain statistical redundancies—which when compressed give way to simple patterns—no such redundancies exist in our task. Thus, we will show, for certain graphs, straightforward notions of succinctness does not break the tie between the two ways of representing the data. Concretely, in generalization theories, the (bit or norm) complexity of a lookup table is larger than the succinct one by a factor that scales polynomially with the size of the training set. In contrast, in the path-star or cycle memorization task, this gap reduces to a constant (that can lie in $[1,2]$ independent of graph size), implying a weak to no implicit capacity pressure:
Proposition 1.
(Geometric and associative memory are roughly equally succinct) For certain graphs, the complexity of the local associative memory both in terms of bits and $\ell_{2}$ norms is either equal to or at most twice that of the global geometric memory (where equality is achieved if there is no weight-tying or if only forward edges are stored [^3] Note that a geometry appears even when storing only one direction of the edges in the smaller graphs where both forms of storage must be equally succinct; see §D.1.2.). (Proof and more explanations in §E.2)

### v2

@@ 2.1 The in-weights path-star task
Task definition. For our study, we consider a task inspired by the above task, yet fundamentally different. In our task, the graph is stored in the model weights rather than presented in context (Fig.˜3). Here, all examples are generated from a (very large) fixed graph $\mathcal{G}$ (of degree $d$ and path length $\ell$), rather than a (small) fresh graph per example. The model is made to memorize the full graph in its weights through “edge-memorization examples”, where the input is some node $v$, and the next-token target is an adjacent node $v^{\prime}$. Separately, path-finding examples are generated from the same graph by picking as input a random leaf node ${v}_{\tt{leaf}}$ (a single token) and as target, the unique root-goal path (a sequence from ${v}_{\tt{root}}$ to ${v}_{\tt{leaf}}$). The model is trained for path-finding on only a subset of such leaves from our fixed graph, and is tested on the remaining “unseen” leaves. By “unseen”, we mean that the paths to these leaves are never seen end-to-end; only the constituent edges are seen individually, which doesn’t trivialize the path-finding task. (Besides this path-star topology, we consider some harder variants, deferred to §C.2.)
@@
Figure 3: Overview of our in-weights path-star task. All examples are derived from a fixed path-star graph. Training involves two types of examples: (i) edge memorization examples; (ii) path-finding examples, where the prefix is some leaf, and the target is the full path. Test examples are path examples corresponding to a held-out set of leaves.
@@
Experimental setup. We use a from-scratch, decoder-only Transformer (GPT-mid) [138]; we corroborate all our findings on Mamba in §C.1. For the most stable results, we interleave the edge-memorization and path-finding examples during training as in Fig.˜3 and also use pause tokens [57]. We found that it is important to provide both forward and reverse edges for edge-memorization in order to dodge the reversal curse [136, 9], but this may not be necessary for smaller graphs; see §D.1. This reverse augmentation is not given for the path-finding examples. Details about the formatting and the hyperparameters are in §B, followed by additional analyses in §D.
2.2 Success of implicit in-weights reasoning
Our first observation is that, despite the adversarially-constructed topology of our graph, and even on path lengths of $6$ to $10$ hops and graphs as large as $10^{4}$ nodes, our next-token-trained models succeed at the in-weights path-finding task. We report this in Figs.˜4 and 14 (left). This isolates an implicit in-weights reasoning sharper than what exists in the literature. Known instances are on small scales of 200 or fewer entities [85, 48] or on 2-hop tasks [174, 43, 189, 74]. We report success in some harder topologies in §C.2.
Next, we isolate an even sharper instance of success. To arrive at that, we ask: why does the next-token learner succeed on massive in-weights path-star tasks, when back in B&N’24 it failed on even tiny in-context path-star tasks? A possible factor could be that in the in-weights setting, the model does not succumb to trivial left-to-right, next-token shortcuts; thus gradients from farther-away tokens survive, allowing the model to discover the true right-to-left solution. However, we find this is likely not the case since the tokens are not learned in a right-to-left order—we report this in Figs.˜4 and 14 (middle) and elaborate on it in §A.1.1. This inspires us to test a harder task where any future-token supervision is manually eliminated: we train the model purely on the first-token loss, conditioned on only the leaf node that is many hops away. Learning the first token—the key decision-making token—in isolation, is (theoretically and empirically) impossible in the in-context path-star task, as B&N’24 show and as reproduced in Fig.˜5 (right). In striking contrast in our in-weights task, our model learns this token in isolation, demonstrating an even sharper instance of implicit reasoning:
@@ 2.2 Success of implicit in-weights reasoning
(Hardest, first token is learned in isolation) On in-weights path-star graphs of as many as $5\times 10^{4}$ nodes, when the model is trained on only edge-memorization examples and first-token-training examples from $75\%$ of all the paths, on unseen paths, both the Transformer and Mamba are able to predict the first token when conditioned on held-out leaves, with as much as 100% accuracy (Figs.˜4 and 14 (right)).
That the hardest (first) token is learned, we argue, is difficult to square with the abstraction that parametric memory strictly stores local associations. We explain this in the next subsection.
Figure 4: Success of Transformer in in-weights path-star task. (§2) (left) A next-token-trained Transformer achieves perfect or highly non-trivial accuracy on large path-star graphs $\mathcal{G}_{d,\ell}$. (middle) Learning order of tokens. The tokens of a path are not learned in the reverse order i.e., the model does not learn the right-to-left solution. Thus, gradients from the future tokens are not critical for success. (right) Success of hardest-token-only task. In fact, the hardest token (the first) given the leaf is learned in isolation to non-trivial accuracy (Section˜2.2). Success of this $\ell$-fold composition task is hard to explain within the associative memory view (§2.3). Analogous plots for Mamba are in Fig.˜14. We invite the reader to contrast these in-weights task results with the in-context task ones in Fig.˜5.
Figure 5: Failure of Transformer in in-context path-star task. (B&N’24) We report the failure of next-token Transformers in the in-context version of the path-star task, reproducing results from B&N’24. (left) Full path accuracy remains at chance level across different small graph sizes. (middle) Learning order of tokens with teacherless (multi-token-trained) objective shows a clear right-to-left learning cascade. (right) The hardest (first) token given the leaf fails to be learned in isolation, contrasting sharply with in-weights success shown in Fig.˜4.
@@ 2.3 The contradiction behind learning the hardest token
Atomic facts are often imagined to be stored in the parameters as local associations, where a matrix acts as a lookup table. We make this concrete. For any tokens $u$ and $v$, let $f(u)[v]$ denote the logit of a sequence model $f$ on the target $v$, conditioned on the input $u$. Then:
(Local associative parametric memory) We say that a deep sequence model $f$ has memorized a graph $\mathcal{G}=({V},E)$ associatively iff for any vertices $u$ and $v$, the logit $f(u)[v]$ is high only on adjacent vertices, $(u,v)\in E$. In its simplest form, this model is typically abstracted as $f(u)[v]={\bf{\Phi}}(u){\bf{W}}_{\tt assoc}{\bf{\Phi}}(v)$. Here, ${\bf{W}}_{\tt assoc}$ encodes the adjacencies as $\sum_{(u,v)\in E}{\bf{\Phi}}(u)\otimes{\bf{\Phi}}(v)$; the embeddings $\bf{\Phi}$} are arbitrary “keys” that by themselves encode no associations in graph $\mathcal{G}$, e.g., they are orthogonal embeddings [177, 51, 78, 202, 27] or random (and thus, near-orthogonal) embeddings [124, 26, 20, 176].
@@
Observe that, in this abstraction, the first node in our task requires composing a local recall function $\ell$-many times: recall the predecessor of ${v}_{\tt{leaf}}$ from the weights, namely $v_{\ell-1},$ then the predecessor of that, $v_{\ell-2}$, and so on. In short, one can write $v_{1}=\texttt{Predec}^{\ell}({v}_{\tt{leaf}})$.
@@
Hardness of learning compositions. With this structure however, the first-token $\ell$-fold composition task should intuitively require $\Omega(\exp(\ell))$ compute to learn—just like the first token was demonstrably hard-to-learn in the in-context path-star task [B&N’24]. One intuition is that the optimization task is a search for a needle in the haystack: there is an $\exp(\ell)$ space of possible discrete compositions, and all but the correct one have an equally miserable loss value; with the loss terrain rendered flat and the gradients uninformative, the learner is forced to sift through the vast hypothesis space to find a needle. As a preliminary corroboration of this intuition, in §E.1, we find that models fail at this task when the embeddings are frozen. We also refer the reader to existing results that empirically demonstrate that compositions are hard to learn [156, 2, 61, 54, 1, 2] and others that theoretically demonstrate hardness [108, 32, 132, 156, 180, 159] (although only in a certain sense; see Remark˜2).
@@
The power of step-wise aid. This barrier could be surmounted if, rather than providing supervision from only the end output of the composition, there was supervision from each hop. This could mean providing a data curriculum [149, 180, 188, 5, 36] or providing chain-of-thought supervision [181, 125] or reducing compositionality through various means. Indeed, prior positive results on implicit in-weights reasoning involve one such aid or the other. First, Khona et al. [85] have paths of varying lengths (see their Fig 11), which provides an implicit curriculum; they also provide full path supervision, not just the first-token loss. Furthermore, their test and train paths overlap, reducing compositionality, potentially allowing the model to stitch substrings it has seen (an ability demonstrated as possible [177]). Other results [174, 43, 189, 74] are on 2-hop tasks, implying a friendly exponent.
Our setting, by design, offers none of these aids. Our paths have fixed lengths (hence, no implicit curriculum), are disjoint (so, no test-train overlaps), and require as much as a $10$-fold composition (so, a daunting exponent). What we isolate, therefore, is a stronger and more sterile instance of implicit in-weights reasoning, one that is cleanly inconsistent within the associative view of parametric memory.
@@ 2.4 Two competing data structures for parametric memory
The paradox begins to resolve with fragmented yet salient observations in literature—[78, 85, 189] and [74]—that we will flesh out more generally. First, [78, 85, 189] find that the embeddings of their tokens reflect some useful notion of distance, aiding their respective implicit reasoning tasks. Independently, [74, 152] find that memorizing associations through an explicitly factorized architecture allows the model to draw unseen (two-hop) inferences. We will generalize these observations to our much larger graphs and many-hop tasks; but first, we point out that these observations are related and lead to profound implications. First, these observations paint an altogether different picture of parametric memory itself: a memory of atomic facts that does not take shape as a lookup over arbitrary embeddings but as a geometry of highly-organized embeddings. While the former directly stores the adjacency matrix $\bf{A}$} in the matrix ${\bf{W}}_{\tt assoc}$ (up to some rotations defined by $\bf{\Phi}$}), the latter “factorizes” the associations into a low-rank form, ${\bf{\Phi}}_{\tt geom}(u)^{T}{\bf{\Phi}}_{\tt geom}(v)$. Importantly, whereas an associative memory only makes local information accessible, a geometric memory readily exposes global multi-hop relationships (§4 discusses what these global relationships are). We define the notion of geometric memory below, and compare the two forms of memory in Table˜1.
@@
Figure 6: Evidence of global geometry of Transformer in path-star task. (a) In the heatmap, entry $(i,j)$ is the cosine distance between the leaf embedding of path $i$ (row) and the first-hop embedding of path $j$ (col). The clear diagonal line implies that embeddings within each path are more aligned, reflecting global structure. (b) UMAP projection of token embeddings where each point is a node embedding; color indicates path identity. Different paths form separated clusters (see Fig.˜27 for clearer image). (c) For a later experiment in §3.1: heatmap for a model trained only on edge-memorization still reveals a level of geometry, although weaker than in (a). Analogous plots for Mamba are in Fig.˜15
(Global geometric parametric memory) We say that a deep sequence model $f$ has memorized a graph $\mathcal{G}=({V},E)$ geometrically iff for any pairs of vertices $u$ and $v$, the logit $f(u)[v]$ reflects some notion of multi-hop closeness in $\mathcal{G}$. At its simplest, this can be abstracted as $f(u)[v]={\bf{\Phi}}_{\tt geom}(u)\cdot{\bf{\Phi}}_{\tt geom}(v)$, where ${\bf{\Phi}}_{\tt geom}$ reflects some graph embedding of the vertices. [^1] In the simple settings of this paper, the geometric embeddings ${\bf{\Phi}}_{\tt geom}$ in practice are found in the token embeddings themselves; but more generally, such an embedding may reside in any other layer. We consider even that to be geometric memorization. [^2] Note that this simplistic abstraction may assign the highest output logit to the input vertex itself (i.e., $\arg\max_{v}{\bf{\Phi}}_{\tt geom}(u)\cdot{\bf{\Phi}}_{\tt geom}(v)=u$) regardless of whether a self-loop $(u,u)$ exists. To account for this, one may incorporate an auxiliary self-loop-detecting circuit. We treat this as a pedantic detail best ignored for a tidy discussion.
@@
The resolution. One can view the two parametric memories of Sections˜2.3 and 2.4 as two competing data structures, both representable by a deep sequence model, each yielding its own learning complexity for the hardest token (much like how a heap or an array would yield different search complexities). The local associative memory is analogous to a linked list: for the first token, this would incur a (benign) $\ell$-hop lookahead step but learning this incurs an exponential cost. The geometric memory is powerful in that it reduces this learning complexity all the way to $\Theta(1)$. For instance, we may expect that the embeddings of all nodes in path $i$ are clustered tightly around a unique path vector $\mathbf{z}_{i}$. In this data structure, learning the first token given the leaf node is a one-hop task: simply find a one-hop neighbor of the central node that is best-aligned with ${\bf{\Phi}}_{\tt geom}({v}_{\tt{leaf}}^{(i)})$. As stated below, such a structure indeed materializes both in our massive graphs and in tiny graphs; this geometry turns out to be fundamental to any gradient-descent-trained deep sequence model and not specific to attention or residual connections:
@@
(Evidence of global geometry across various deep sequence models) For our large path-star graphs, the (token) embeddings of the leaf and first node of a path are clustered closer to each other than those of other paths in both the Transformer and Mamba (Fig.˜6 and Fig.˜15). Similar geometries arise for a variety of small graphs for the Transformer, Mamba and a simple 3-layer neural network model without attention and residual connections—see visualizations in Fig.˜1 and Section˜C.3; §C.3: Figs.˜18, 19, 20 and 21) and node-node cosine-similarity heatmaps (§C.3; Figs.˜24 and 23).
@@
While these geometries help make sense of why implicit reasoning succeeds, it also leaves us with a foundational question of why the model prefers one form of memory over the other. We elaborate on this in the next section.
3 The memorization puzzle: The emergence of geometric memory is not easy to explain
During training, the two types of parametric memory—two equally valid ways to fit the training data—must compete with each other. Why does the geometric memory prevail over the associative? We demonstrate that this behavior is hard to explain.
The reader may offer various plausible explanations for this behavior by borrowing intuition from other well-known phenomena. Some readers may find the rise of a geometry familiar in light of well-known geometries of features in language [40, 130, 112] and arithmetic tasks [100, 120, 117, 59]. Some others may find the phenomenon reminiscent of the competition between memorization and generalization studied in the generalization puzzle [195, 122, 17, 23] in regression and classification settings. These settings suggest that representations arise due to pressures from either the supervision or the architecture or the optimizer; such pressures forbid a lookup. Thus, it may be tempting to cite these well-known learning pressures as plausible explanations for what we observe in our sequence models. However, we design experiments refuting such explanations: we show that a geometric representation arises over the associative lookup even in the absence of such learning pressures. Through this analysis, we pose the memorization puzzle in sequence modeling.